OpenStudy (anonymous):

BROS COME HELP..The radius of the Earth is approximately 3,960 miles. Find the approximate surface-area-to-volume ratio of the Earth. A. 0.00025 B. 0.00076 C. 1,320 D. 11,880

OpenStudy (anonymous):

The radius of the Earth is approximately 3,960 miles. Find the approximate surface-area-to-volume ratio of the Earth. A. 0.00025 B. 0.00076 C. 1,320 D. 11,880

OpenStudy (anonymous):

The volume of a sphere is \[\frac{ 4 }{ 3 } \pi r ^{3}\]

OpenStudy (anonymous):

The surface area of a sphere is \[4 \pi r ^{2}\]

OpenStudy (anonymous):

Divide the surface area by the volume to get \[\frac{ 1 }{ \frac{ 1 }{ 3 } r}\]

OpenStudy (anonymous):

Plug in r = 3,960 miles so Surface/Volume = .00076 which corresponds to choice B.